OpenStudy (anonymous):

solve the following proportion for x: 2/11=x/91-x a.70.78 b.14 c.20.22 d.77

OpenStudy (anonymous):

cross multiply: 11x = 2(91-x) 11x = 182-2x 11x+2x =182 13x=182 x=??

OpenStudy (anonymous):

14

OpenStudy (anonymous):

thanks